A particular reaction has a ΔHo value of -184 kJ and ΔGo of -124 kJ at 217 K. Calculate ΔSo at 217 K in J/K
1
Expert's answer
2019-10-28T07:49:50-0400

We have to use a so-called Gibbs equation:


"\delta G^0=\delta H^0-T \delta S^0;"


From this, we will get the value of entropy:


"\delta S^0=\frac{\delta H^0-\delta G^0}{T}=\frac{-184kJ-(-124kJ)}{217K}=-0.276kJ/K(-276J/K)."


Thus, the entropy of this reaction would be -276 J/K.